# Fundamental Limits of Coded Polynomial Aggregation

### v1

@@ Abstract
We introduce the problem of coded polynomial aggregation (CPA) in distributed computing systems, where the goal is to recover a weighted sum of polynomial evaluations using responses from distributed workers. A fundamental question in CPA is the minimum number of responses required to recover the computation. A straightforward approach is individual decoding, where the system first decodes all individual polynomial evaluations and then aggregates them, which requires at least d(K1)+1d(K-1)+1 responses for a degree-$d$ polynomial evaluated on $K$ data points. In this paper, we establish a necessary and sufficient condition for exact recovery when the number of responses satisfies $N\leq d(K-1)$, a regime where individual decoding is infeasible. Moreover, we characterize the minimum number of responses required by CPA, showing that it equals $\left\lfloor\frac{K-1}{2}\right\rfloor+1$ when $d=1$, and (d1)(K1)+1(d-1)(K-1)+1 when $d\geq 2$. These thresholds are strictly smaller than those required by a straightforward strategy based on individual decoding. We further provide explicit constructions of CPA schemes that achieve this bound.
@@ I Introduction
Coded distributed computing is a fundamental paradigm for executing large-scale data processing tasks, where a master node partitions a computation task, assigns sub-tasks to multiple workers, and aggregates their responses to recover the desired computation. One commonly used metric in coded distributed computing systems is the minimum amount of information that must be collected from the workers to recover the computation. This metric can be characterized by the minimum number of responses, defined as the minimum number of worker responses required for successful decoding.
@@
For matrix multiplications and polynomial computations, many prior works apply polynomial-based coded computing techniques. Specifically, the authors in [5] use one-dimensional maximum distance separable (MDS) codes to encode matrices, achieving a minimum number of responses equal to the partition level of the data matrix. Product codes [6] use two-dimensional MDS codes and apply iterative decoding algorithms. Polynomial codes [14] encode sub-matrices as coefficients of a univariate polynomial and interpolate this polynomial using responses from workers. PolyDot codes and MatDot codes [1] extend the interpolation-based approach to trade off computation and communication costs by partitioning datasets. Entangled polynomial codes [12, 15] exploit bilinear structure and extend polynomial coded computing to scenarios such as secure, private, and batch computation. To construct numerically stable interpolation schemes, the authors in [2] apply Chebyshev polynomials to coded computing. Instead of univariate polynomials, bivariate polynomial coded schemes [3] develop two-dimensional interpolation, enabling finer trade-offs among storage, computation, and communication. Differing from encoding sub-computations as coefficients of a polynomial, Lagrange coded computing [13] encodes sub-computations as evaluations of Lagrange polynomials, achieving lower decoding complexity.
Apart from decoding exact computations, several works consider approximate computation of arbitrary functions. The Berrut-approximated coded computing scheme proposed in [4] replaces polynomial interpolation with rational interpolation to enable numerically stable approximation. In 2025, the authors in [8] proposed learning-theoretic coded computing, by formulating coded computing as an encoder–decoder design problem that minimizes the mean squared error (MSE), and derived optimal coding schemes based on smoothing splines. This approach is extended to the scenario where each worker becomes a straggler with some probability [7].
@@
A common paradigm underlying many existing coded computing schemes is individual decoding, where the master reconstructs all individual sub-computations before computing the desired output. In particular, for matrix multiplications, the works in [5, 6, 14, 1, 12, 15, 2, 3] decode all sub-matrices contained in the desired matrix, and therefore recover the computation. However, for aggregation tasks, individual decoding is sufficient but not necessary, and may lead to redundant computation. To compute aggregation tasks in the form of gradient computations, [11] proposed gradient coding, which introduces redundancy into gradient computations to mitigate slow or unavailable workers. A different class of aggregation schemes considers linearly separable computations [10], where the authors focus on computing weighted aggregations of arbitrary computation functions while minimizing communication under uncoded data storage.
@@
Despite these extensive efforts, the problem of computing a weighted aggregation of polynomial computations has not been studied as a standalone objective. Specifically, we consider the computation of weighted aggregation,
@@
where $w_{k}\in\mathbb{C}$ are non-zero weights, and $F(\cdot)$ is a polynomial of degree $d$ that operates element-wise on each data matrix $\bm{X}_{k}\in\mathbb{C}^{q\times v}$. This formulation commonly arises in distributed learning and signal processing, where the objective is to recover an aggregated output rather than all individual computations. A natural baseline in coded polynomial computing is individual decoding, where workers compute polynomial evaluations on encoded data, and the master interpolates the resulting polynomial to recover all $F(\bm{X}_{k})$ for $k\in[K]$. Under individual decoding, the required number of responses for exact recovery is d(K1)+1d(K-1)+1. However, the objective is to recover a weighted aggregation rather than each $F(\bm{X}_{k})$, which leads to the following questions.
@@
    We propose a framework of CPA to compute the weighted aggregation in (1), and establish a necessary and sufficient condition for exact recovery in the regime where individual decoding is infeasible.
@@ II Problem Formulation
We consider a distributed computing system consisting of a master node and a set of $N$ worker nodes, indexed by $N]\triangleq\{0,1,\ldots,N-1\}$. Given $K$ data matrices $\bm{X}_{k}\in\mathbb{C}^{q\times v}$ for $k\in[K]$, a polynomial function $F(\cdot)$ of degree $d$, and a weight vector $\bm{w}\in\mathbb{C}^{K}$, where each element $w_{k}\neq 0$ for all $k\in[K]$, the objective of the master is to compute the weighted aggregation defined in (1), using the responses from workers. Each worker returns one response to the master.
@@ II-A1 Encoding
The master selects a set of $K$ distinct numbers $\{\alpha_{k}$ $\in$ $\mathbb{C}:$ $k\in[K]\}$, referred to as the data points, and interpolates an encoder polynomial $E(z)$ such that $E(\alpha_{k})$ == $\bm{X}_{k}$ for all $k\in[K]$. Next, the master selects another set of $N$ distinct numbers $\{\beta_{n}\in\mathbb{C}:$ $n\in[N]\}$, referred to as the evaluation points, satisfying $\{\alpha_{k}:$ $k\in[K]\}$ $\cap$ $\{\beta_{n}:$ $n\in[N]\}$ == $\emptyset$. The master evaluates $E(z)$ at $\{\beta_{n}:n\in[N]\}$ and sends the coded matrix $E(\beta_{n})$ to worker $n$.
@@ II-A3 Decoding
Upon receiving the responses from all $N$ workers, the master interpolates a decoder polynomial $D(z)$ such that $D(\beta_{n})=F(E(\beta_{n}))$ for all $n\in[N]$. The master computes $D(\alpha_{k})$ for $k\in[K]$, and obtains
$\widehat{\bm{Y}}\triangleq\sum_{k=0}^{K-1}w_{k}\,D(\alpha_{k}).$
@@
Fix positive integers $K$, $d$, and $N$, data points $\{\alpha_{k}$ :: $k\in[K]\}$ and evaluation points $\{\beta_{n}:n\in[N]\}$. A CPA scheme is feasible if it produces $\widehat{\bm{Y}}=\bm{Y}$. $\diamond$
@@
For positive integers $K$ and $d$, and a given set of data points $\{\alpha_{k}:k\in[K]\}$, the minimum number of responses is defined as the smallest integer $N$ for which there exist evaluation points $\{\beta_{n}:n\in[N]\}$ such that a feasible CPA scheme exists. $\diamond$
In this paper, the goal is to characterize the minimum number of responses for which a feasible CPA scheme exists, and to construct explicit schemes that achieve this bound.
@@ II-B Individual Decoding Baseline
Existing results on polynomial coded computing, including [14, 1, 12, 15, 2, 3, 13], recover the desired computation by decoding all individual sub-computations through reconstructing polynomial evaluations or coefficients. Among these works, we set Lagrange coded computing [13] as the baseline to our setting. It focuses on polynomial computations and encodes each data matrix $\bm{X}_{k}$ as an evaluation of an encoder polynomial $E(z)$. When applied to the CPA setting, Lagrange coded computing leads to the following CPA based on individual decoding.
@@
Figure 1: Feasible regimes of CPA and the baseline CPA scheme based on individual decoding for computation degrees $d=1$ and $d=2$. For a fixed $K$, the feasible regime consists of the values of $N$ for which the corresponding scheme is feasible. The red region corresponds to CPA, which enables feasibility with fewer responses ($N\leq d(K-1)$) via the orthogonality conditions in Theorem 1, with the minimum number of responses $N^{\star}$ characterized in Theorem 2. The gray region corresponds to the baseline scheme based on individual decoding, which is feasible only when $N\geq d(K-1)+1$, as stated in Lemma 1.
The following lemma characterizes the minimum number of responses for a CPA scheme based on individual decoding.
@@
Proof:
The master recovers all individual evaluations by interpolating the composed polynomial $D(z)=F(E(z))$, so that $D(\alpha_{k})=F(\bm{X}_{k})=F(E(\alpha_{k}))$ for all $k\in[K]$. Since $\deg(E)\leq K-1$ and $F(z)$ has degree $d$, we have $\deg(D)\leq d(K-1)$. Therefore, $D(z)$ can be uniquely interpolated from the worker responses if and only if $N\geq d(K-1)+1$. ∎
Lemma 1 shows that CPA schemes based on individual decoding require at least d(K1)+1d(K-1)+1 responses.
In this paper, we study CPA in the regime $N\leq d(K-1)$, where individual decoding is infeasible. We show that exact recovery remains possible with strictly fewer worker responses by directly exploiting the aggregation structure.
@@ III-B Minimum Number of Responses
The following theorem characterizes the minimum number of responses required for the existence of a feasible CPA scheme for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, where the data points are chosen outside a proper algebraic variety depending on the system parameters $K$, NN, and $d$, and satisfy $\alpha_{i}\neq\alpha_{j}$ for all $i\neq j$.
@@
For positive integers $K$, $d$, and $N$ with $2\leq N\leq d(K-1)$, and for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, there exists evaluation points $\{\beta_{n}:n\in[N]\}$ such that CPA is feasible if and only if $N\geq N^{*}$, where
@@
From Theorem 2, when $d=1$, the minimum number of responses required by CPA scales to approximately half of that required by individual decoding as $K$ grows large. When $d\geq 2$, CPA reduces the required number of responses by K1K-1 compared to the individual decoding baseline in Lemma 1.
Fig. 1 illustrates the minimum number of responses required by CPA and the baseline CPA schemes based on individual decoding. As shown in Fig. 1, for both $d=1$ and $d=2$, CPA has a strictly smaller minimum number of responses for any $K$. Moreover, the difference between the proposed CPA schemes and the baseline CPA scheme based on individual decoding grows with $K$, indicating that the benefit of CPA becomes more pronounced as the number of data matrices increases.
III-C Achievable CPA Schemes
Theorem 3
Proof:
The proof is provided in Section V. ∎
@@ III-C Achievable CPA Schemes
7:Form the polynomial $P_{\text{form}}(z)=\sum_{n=0}^{N}c_{n}z^{n}$
@@
In contrast, CPA allows $N=2$, which gives $C=1$. Applying Algorithm 1, we obtain $\bm{V}=\begin{bmatrix}1&1&1\end{bmatrix}\text{ and }\bm{A}=\begin{bmatrix}1&-1&1\\ 1&0&0\\ 1&1&1\end{bmatrix}$, which yields $\bm{U}=\begin{bmatrix}1&1&0\end{bmatrix}$. Solving $\bm{U}\bm{c}=\bm{0}$ gives $\bm{c}$ =(0.0506=(-0.0506, $0.0506$, $0.5)$ and $P_{\text{form}}(z)$ =0.0506=-0.0506 +0.0506z+0.0506z $0.5z^{2}$. The two roots are $\beta_{0}=-0.37272$ and $\beta_{1}=0.27152$, which are distinct and disjoint from $\{-1,0,1\}$. Finally, the orthogonality conditions in Theorem 1 are numerically satisfied, with $\sum_{k=0}^{2}w_{k}P(\alpha_{k})\approx 5.5\times 10^{-17}$.
@@ IV Proof of Theorem 1
We start by considering the scalar case where $\bm{X}_{k}$ reduces to a scalar $x_{k}$ and the desired output $\bm{Y}$ reduces to a scalar $y$, i.e., $q=v=1$. The extension to the matrix case follows directly by applying the same argument element-wise, since the polynomial $F(\cdot)$ operates element-wise on the data matrices. Define the error polynomial $\Delta(z)\triangleq D(z)-F(E(z))$. Its degree satisfies $\deg(\Delta)\leq\max\{\deg(D),\deg(F(E))\}\leq\max\{N-1,d(K-1)\}=d(K-1)$. The decoding error $\widehat{y}-y$ can be expressed by
@@
where $P(z)=\prod_{n\in[N]}(z-\beta_{n})$ has degree $N$, as shown in (3), and $R(z)$ is a polynomial satisfying $\deg(R)$ == $\deg(\Delta)-\deg(P)$ $\leq$ d(K1)Nd(K-1)-N. Writing $R(z)$ == $\sum_{j=0}^{\deg(R)}r_{j}z^{j}$, where $\{r_{j}:$ $j\in[\deg(R)+1]\}$ are arbitrary complex coefficients. We substitute (6) into (5) yields
$\widehat{y}-y$ $=\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})R(\alpha_{k})$
$=\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\sum_{j=0}^{\deg(R)}r_{j}\alpha_{k}^{j}$
$=\sum_{j=0}^{\deg(R)}r_{j}\left(\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}\right).$
Therefore, $\widehat{y}-y=0$ for all admissible choices of $R(z)$ if and only if $\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}=0$ for all $j=$ 0, 11, $\ldots$, d(K1)Nd(K-1)-N. This establishes the necessary and sufficient condition for feasibility in the scalar case. The same argument applies element-wise to the matrix-valued case, and hence the condition remains necessary and sufficient for feasibility. This completes the proof of Theorem 1.
V Proof of Theorem 2 and Theorem 3

### v2

@@ Abstract
We introduce the problem of coded polynomial aggregation (CPA) in distributed computing systems, where the goal is to recover a weighted sum of polynomial computations using the responses from distributed workers. A fundamental question in CPA is the minimum number of responses required to exactly recover the computation. A straightforward strategy is individual decoding, where the system first decodes all individual polynomial computations and then aggregates them, which requires at least d(K1)+1d(K-1)+1 responses for a degree-$d$ polynomial computed on $K$ datasets. In this paper, we study CPA in the regime where the number of responses satisfies $N\leq d(K-1)$, and show that exact recovery can be achieved by directly exploiting the aggregation structure and the algebraic properties of polynomials. In particular, we establish a necessary and sufficient condition for exact recovery in CPA. Moreover, we characterize the minimum number of responses, showing that it equals $\left\lfloor\frac{K-1}{2}\right\rfloor+1$ when $d=1$, and (d1)(K1)+1(d-1)(K-1)+1 when $d\geq 2$. These thresholds are strictly smaller than that required by individual decoding. We further provide explicit constructions of CPA schemes that achieve this bound.
@@ I Introduction
Coded distributed computing is a fundamental paradigm for executing large-scale data processing tasks, where a master node partitions a computation task, assigns sub-tasks to multiple workers, and aggregates their responses to recover the desired computation. One commonly used metric in coded distributed computing systems is the minimum amount of information that must be collected from the workers to recover the computation. This metric can be characterized by the minimum number of responses, i.e., the smallest number of worker responses required for exact recovery.
@@
For matrix multiplications and polynomial computations, many prior works apply polynomial-based coded computing techniques. Specifically, the authors in [5] use one-dimensional maximum distance separable (MDS) codes to encode matrices, achieving a minimum number of responses, which equals the partition level of the data matrix. Product codes [6] use two-dimensional MDS codes and apply iterative decoding algorithms. Polynomial codes [14] encode sub-matrices as coefficients of a univariate polynomial and interpolate this polynomial using responses from workers. PolyDot codes and MatDot codes [1] extend the interpolation-based approach to trade off computation and communication costs by partitioning datasets. Entangled polynomial codes [12, 15] exploit a bilinear structure and extend polynomials codes to scenarios such as secure, private, and batch computation. To construct numerically stable interpolation schemes, the authors in [2] apply Chebyshev polynomials to coded computing. Instead of univariate polynomials, bivariate polynomial coded schemes [3] develop two-dimensional interpolation, enabling finer trade-offs among storage, computation, and communication. Differing from encoding sub-computations as coefficients of a polynomial, Lagrange coded computing [13] encodes sub-computations as evaluations of Lagrange polynomials, achieving a lower decoding complexity. Apart from decoding exact computations, several works consider approximated coded computing for arbitrary functions under numerical or probabilistic guarantees [4, 8, 7].
@@
A common paradigm underlying many existing coded computing schemes is individual decoding, where the master reconstructs all individual sub-computations before computing the desired output. In particular, for matrix multiplications, the works in [5, 6, 14, 1, 12, 15, 2, 3] decode all sub-matrices contained in the desired matrix, and therefore recover the computation. However, for aggregation tasks, individual decoding is sufficient for exact recovery but may lead to redundant computations. To compute aggregation tasks in the form of gradient computations, [11] proposed gradient coding, which introduces redundancy into gradient computations to mitigate slow workers. A different class of aggregation schemes considers linearly separable computations [10], where the authors focus on computing weighted aggregations of arbitrary computation functions with minimizing the communication cost.
@@
Despite these extensive efforts, the problem of computing a weighted aggregation of polynomial computations has not been studied as a standalone objective. Specifically, we consider the computation of a weighted aggregation,
@@
where $w_{k}\in\mathbb{C}$ are non-zero weights, and $F(\cdot)$ is a polynomial of degree $d$ that operates element-wise on each data matrix $\bm{X}_{k}\in\mathbb{C}^{q\times v}$. This formulation commonly arises in distributed learning and signal processing, where the objective is to recover an aggregated output rather than all individual computations. A straightforward strategy is individual decoding, where workers compute polynomial computations on encoded data, and the master interpolates the resulting polynomial to recover all $F(\bm{X}_{k})$ for $k\in[K]$. Under individual decoding, the minimum number of responses for exact recovery is d(K1)+1d(K-1)+1. However, the objective is to recover a weighted aggregation rather than each $F(\bm{X}_{k})$, which leads to the following questions.
@@
    We propose a CPA framework for computing the weighted aggregation in (1), and establish a necessary and sufficient condition for exact recovery in the regime $N\leq d(K-1)$, where individual decoding is infeasible.
@@ II Problem Formulation
We consider a distributed computing system consisting of a master node and a set of $N$ worker nodes, indexed by $N]\triangleq\{0,1,\ldots,N-1\}$. Given $K$ data matrices $\bm{X}_{k}\in\mathbb{C}^{q\times v}$ for $k\in[K]$, a polynomial function $F(\cdot)$ of degree $d$ that operates element-wise on each data matrix, and a weight vector $\bm{w}\in\mathbb{C}^{K}$, where each element $w_{k}\neq 0$ for $k\in[K]$, the objective of the system is to compute the weighted aggregation defined in (1), using the responses from workers. Each worker returns one response to the master.
@@ II-A1 Encoding
The master selects a set of $K$ distinct data points $\{\alpha_{k}\in\mathbb{C}:k\in[K]\}$, and interpolates an encoder polynomial $E(z)$ such that $E(\alpha_{k})=$ $\bm{X}_{k}$ for all $k\in[K]$. Next, the master selects a set of $N$ distinct evaluation points $\{\beta_{n}$ $\in\mathbb{C}:$ $n\in[N]\}$ satisfying $\{\alpha_{k}:$ $k\in[K]\}$ $\cap$ $\{\beta_{n}:$ $n\in[N]\}$ == $\emptyset$. The master evaluates $E(z)$ at $\{\beta_{n}:n\in[N]\}$ and sends the coded matrix $E(\beta_{n})$ to worker $n$.
@@ II-A3 Decoding
Upon receiving responses from all $N$ workers, the master interpolates a decoder polynomial $D(z)$ such that $D(\beta_{n})=F(E(\beta_{n}))$ for all $n\in[N]$. The master then evaluates $D(z)$ at the data points $\{\alpha_{k}:k\in[K]\}$ and obtains
$\widehat{\bm{Y}}\triangleq\sum_{k=0}^{K-1}w_{k}\,D(\alpha_{k}).$ (2)
@@
Fix positive integers $K$, $d$, and $N$, data points $\{\alpha_{k}$ :: $k\in[K]\}$ and evaluation points $\{\beta_{n}:n\in[N]\}$. A CPA scheme is feasible if $\widehat{\bm{Y}}=\bm{Y}$. $\diamond$
In this paper, we treat the data points $\{\alpha_{k}:k\in[K]\}$ as fixed system parameters [^1] Allowing joint design of the data points $\{\alpha_{k}:k\in[K]\}$ and the evaluation points $\{\beta_{n}:n\in[N]\}$ may further enlarge the feasible design space of CPA schemes.. We assume that they are pairwise distinct, i.e., $\alpha_{i}\neq\alpha_{j}$ for all $i\neq j$, and generic [^2] For background on relevant concepts in algebraic geometry, such as genericity, algebraic varieties, Lebesgue measure zero, and Zariski-open dense subsets used in this paper, we refer the reader to [9].in the sense that they are chosen outside a proper algebraic variety determined by the system parameters $K$, NN, $d$ and $\bm{w}$.
@@
For positive integers $K$ and $d$, and a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, the minimum number of responses is defined as the smallest integer $N$ such that there exists a feasible CPA scheme. $\diamond$
In this paper, the goal is to characterize the minimum number of responses and to construct explicit schemes that achieve this bound.
@@ II-B Individual Decoding Baseline
Existing results on polynomial codes, including [14, 1, 12, 15, 2, 3, 13], recover the desired computation by decoding all individual sub-computations through reconstructing polynomial evaluations or coefficients. Among these works, we set Lagrange coded computing [13] as the baseline to our setting. It focuses on polynomial computations and encodes each data matrix $\bm{X}_{k}$ as an evaluation of an encoder polynomial $E(z)$. When applied to the CPA setting, Lagrange coded computing leads to the following CPA based on individual decoding.
@@
Figure 1: Feasible regimes of CPA and the baseline CPA scheme based on individual decoding for computation degrees $d=1$ and $d=2$. For a fixed $K$, the feasible regime consists of the values of $N$ for which the corresponding scheme is feasible. The red region corresponds to CPA, which enables feasibility with fewer responses ($N\leq d(K-1)$) via the orthogonality conditions in Theorem 1, with the minimum number of responses $N^{*}$ characterized in Theorem 2. The gray region corresponds to the baseline scheme based on individual decoding, which is feasible only when $N\geq d(K-1)+1$, as stated in Lemma 1.
From standard polynomial interpolation arguments, the following lemma characterizes the minimum number of responses for CPA based on individual decoding.
@@
In this paper, we study CPA in the regime $N\leq d(K-1)$, where a CPA scheme based on individual decoding is infeasible. We show that exact recovery is achievable by directly exploiting the aggregation structure and the algebraic properties of polynomials.
@@ III-B Minimum Number of Responses
The minimum number of responses of CPA is characterized as follows.
@@
For positive integers $K$, $d$, and $N$ with $2\leq N\leq d(K-1)$, and for a given generic set of pairwise distinct data points $\{\alpha_{k}:k\in[K]\}$, there exist evaluation points $\{\beta_{n}:n\in[N]\}$ such that CPA is feasible if and only if $N\geq N^{*}$, where
@@
From Theorem 2, when $d=1$, the minimum number of responses of CPA scales to approximately half of that required by individual decoding as $K$ grows large. When $d\geq 2$, CPA reduces the required number of responses by K1K-1 compared to the individual decoding baseline in Lemma 1.
Fig. 1 illustrates the feasible regimes and the minimum number of responses, required by CPA and the baseline CPA scheme based on individual decoding. From Fig. 1, for both $d=1$ and $d=2$, CPA has a strictly smaller minimum number of responses for any $K$. Moreover, the difference between the proposed CPA schemes and the baseline CPA scheme based on individual decoding grows with $K$, indicating that the benefit of CPA becomes more pronounced as the number of data matrices increases.
@@
7:Construct the polynomial $P_{\text{form}}(z)=\sum_{n=0}^{N}c_{n}z^{n}$
@@
III-C Achievable CPA Schemes
Theorem 3
Proof:
The correctness of Algorithm 1 follows directly from the constructive sufficiency proof of Theorem 2. Importantly, all but a Lebesgue measure-zero subset of vectors $\bm{c}\in\ker(\bm{U})$ lead to polynomials $P_{\text{form}}(z)$ such that its roots $\{\beta_{n}:n\in[N]\}$ are distinct and satisfy the orthogonality conditions in (3) and $\{\beta_{n}:n\in[N]\}\cap\{\alpha_{k}:k\in[K]\}=\emptyset$. Details are provided in Section V-C. ∎
@@ III-C Achievable CPA Schemes
In contrast, CPA allows $N=2$, which gives $C=1$. Applying Algorithm 1, we obtain $\bm{V}=\begin{bmatrix}1&1&1\end{bmatrix}\text{ and }\bm{A}=\begin{bmatrix}1&-1&1\\ 1&0&0\\ 1&1&1\end{bmatrix}$, which yields $\bm{U}=\begin{bmatrix}1&1&0\end{bmatrix}$. Solving $\bm{U}\bm{c}=\bm{0}$ gives $\bm{c}$ =(0.0506=(-0.0506, $0.0506$, $0.5)$ and $P_{\text{form}}(z)$ =0.0506=-0.0506 +0.0506z+0.0506z $0.5z^{2}$. The two roots are $\beta_{0}=-0.37272$ and $\beta_{1}=0.27152$, which are distinct and disjoint from $\{-1,0,1\}$. Finally, the orthogonality conditions in Theorem 1 are numerically satisfied, with $\sum_{k=0}^{2}w_{k}P(\alpha_{k})\alpha_{k}^{0}\approx 5.5\times 10^{-17}$.
@@ IV Proof of Theorem 1
We first consider the scalar case where $\bm{X}_{k}$ reduces to a scalar $x_{k}$ and the desired output $\bm{Y}$ reduces to $y$. The extension to the matrix case follows directly by applying the same argument element-wise, since the polynomial $F(\cdot)$ operates element-wise on the data matrices. Define the error polynomial $\Delta(z)\triangleq D(z)-F(E(z))$, which satisfies $\deg(\Delta)$ $\leq$ $\max\{\deg(D)$, $\deg(F(E))\}$ $\leq$ $\max\{N-1$, d(K1)}=d(K1)d(K-1)\}=d(K-1). The recovery error $\widehat{y}-y$ can be expressed by
@@
where $P(z)=\prod_{n\in[N]}(z-\beta_{n})$ has degree $N$, as shown in (4), and $R(z)$ is a polynomial satisfying $\deg(R)$ == $\deg(\Delta)-\deg(P)$ $\leq$ d(K1)Nd(K-1)-N. Writing $R(z)$ == $\sum_{j=0}^{\deg(R)}r_{j}z^{j}$, where $\{r_{j}:$ $j\in[\deg(R)+1]\}$ are arbitrary complex coefficients. Substituting (7) into (6) yields
$\widehat{y}-y=\sum_{j=0}^{\deg(R)}r_{j}\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j},$
as $\widehat{y}-y$ == $\sum_{k=0}^{K-1}$ $w_{k}P(\alpha_{k})R(\alpha_{k})$ == $\sum_{k=0}^{K-1}$ $w_{k}P(\alpha_{k})$ $\sum_{j=0}^{\deg(R)}r_{j}\alpha_{k}^{j}$ $\sum_{j=0}^{\deg(R)}r_{j}(\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j})$. Therefore, $\widehat{y}-y=0$ for all admissible choices of $R(z)$ if and only if $\sum_{k=0}^{K-1}w_{k}P(\alpha_{k})\alpha_{k}^{j}=0$ for all $j=$ 0, 11, $\ldots$, d(K1)Nd(K-1)-N. This establishes the necessary and sufficient condition for feasibility in the scalar case. The same argument applies element-wise to the matrix-valued case, and hence the condition remains necessary and sufficient for feasibility. This completes the proof of Theorem 1.
V Proof of Theorem 2